# An Elementary Approach to Scheduling in Generative Diffusion Models

### v1

@@ Abstract
An elementary approach to characterizing the impact of noise scheduling and time discretization in generative diffusion models is developed. Considering a simplified model where the source distribution is multivariate Gaussian with a given covariance matrix, the explicit closed-form evolution trajectory of the distributions across reverse sampling steps is derived, and consequently, the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output is obtained. The effect of the number of time discretization steps on the convergence of this KL divergence is studied via the Euler-Maclaurin expansion. An optimization problem is formulated, and its solution noise schedule is obtained via calculus of variations, shown to follow a tangent law whose coefficient is determined by the eigenvalues of the source covariance matrix. For an alternative scenario, more realistic in practice, where pretrained models have been obtained for some given noise schedules, the KL divergence also provides a measure to compare different time discretization strategies in reverse sampling. Experiments across different datasets and pretrained models demonstrate that the time discretization strategy selected by our approach consistently outperforms baseline and search-based strategies, particularly when the budget on the number of function evaluations is very tight.
@@ I Introduction
Aiming at understanding from first principles, the interplay between noise scheduling and sampling efficiency, in this work, we adopt an elementary approach based on an idealized case where the source distribution is multivariate Gaussian with a given covariance matrix. Such a simple model induces a linear form of the optimal posterior estimator, and consequently an explicit closed-form characterization of the evolution trajectory of the distributions across reverse sampling steps. Using this closed-form formulation, we can conveniently investigate the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output.
@@
Notation: For brevity, we abbreviate a function $f(t)$ as $f$ and denote its discrete-time counterpart $f(t_{j})$ by $f_{t_{j}}$. Derivatives with respect to $t$ are written as $\dot{f}=\mathrm{d}f/\mathrm{d}t$ and $\ddot{f}=\mathrm{d}^{2}f/\mathrm{d}t^{2}$. Throughout $\log(\cdot)$ represents the natural logarithm.
@@ II Preliminaries
In this section, we first review the mechanism of DMs, encompassing the forward diffusion process with its noise schedule design, and the deterministic reverse sampling process employed for generation. Subsequently, we outline the multivariate Gaussian setting that enables a tractable theoretical analysis.
II-A Diffusion Models
@@
where $t\in[0,1]$ denotes the diffusion time, $\mathbf{x}_{0}\sim q(\mathbf{x})$ represents the source data sample, and $\alpha_{t},\sigma_{t}$ parameterize the noise schedule that directly controls the noise-adding process.
Commonly adopted noise schedules include two types: variance-preserving (VP) and variance-exploding (VE) [20]. The VP setting usually satisfies the constraint $\alpha^{2}(t)+\sigma^{2}(t)=1$, where $\alpha(t)$ decays from 1 to 0. In contrast, the VE setting maintains $\alpha(t)\equiv 1$, while $\sigma(t)$ increases from 0 to a predefined maximum $\sigma_{\max}\gg 1$. In either case, the forward transition kernel is given by
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}\!\left(\alpha_{t}\mathbf{x}_{0},\,\sigma^{2}_{t}\mathbf{I}\right),$ (2)
which progressively diffuses the source distribution $q(\mathbf{x}_{0})$ towards a Gaussian noise as $t\to 1$.
The deterministic reverse sampling process, which reverses the forward diffusion process (1) to generate data samples from $q(\mathbf{x})$, is governed by the following update formula [19, 12], for each step $j\in\left[1,N\right]$:
where $\{t_{i}\}_{i=0}^{N}$ denotes the time discretization sequence decreasing from $t_{N}=1$ to $t_{0}=0$, and $\bm{\epsilon}_{\theta}$ is a learned noise estimator. The reverse sampling process is initialized with $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for the VP setting and $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{\max}^{2}\mathbf{I})$ for the VE setting.
To obtain the noise estimator $\bm{\epsilon}_{\theta}$, the training objective is to minimize the following mean squared error [5]:
@@
$\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}\left[\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}}\bigg|\mathbf{x}_{t}\right]=\frac{\mathbf{x}_{t}-\alpha_{t}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\sigma_{t}}.$ (5)
@@
Thus, learning to predict the noise $\bm{\epsilon}$ is mathematically equivalent to estimating the posterior mean of the source data $\mathbf{x}_{0}$.
II-B Gaussian Setup
To facilitate theoretical analysis, we consider the idealized case where the source distribution is multivariate Gaussian with a given covariance matrix, i.e., $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{\mathbf{x}})$. Without loss of generality, we assume that the covariance matrix $\bm{\Sigma}_{\mathbf{x}}$ is positive definite. Let the eigendecomposition of the covariance matrix be $\bm{\Sigma}_{\mathbf{x}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{-1}$, where $\mathbf{U}$ is an orthogonal matrix and $\mathbf{\Lambda}=\text{diag}(\mu_{1},\ldots,\mu_{k})$ denotes the diagonal matrix of eigenvalues. Furthermore, due to the linearity of the forward diffusion process (1), the marginal distribution $q(\mathbf{x}_{t})$ remains Gaussian:
@@ II-B Gaussian Setup
$q(\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{0},\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right).$ (6)
@@
Moreover, the optimal posterior estimator admits a closed-form solution [15]:
@@
III KL-Based Analysis and Applications
In this section, we present the main results derived from our theoretical analysis under the Gaussian setting, and demonstrate their applications. For clarity of exposition, we focus on the VP setting; analogous results for the VE setting are briefly described in discussions and remarks.
III-A KL Divergence Analysis

### v2

@@ Abstract
An elementary approach to characterizing the impact of noise scheduling and time discretization in generative diffusion models is developed. We first utilize the Cramér-Rao bound to identify the Gaussian setting as a fundamental performance limit, necessitating its study as a reference. Building on this insight, we consider a simplified model in which the source distribution is a multivariate Gaussian with a given covariance matrix, together with the deterministic reverse sampling process. The explicit closed-form evolution trajectory of the distributions across reverse sampling steps is derived, and consequently, the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output is obtained. The effect of the number of time discretization steps on the convergence of this KL divergence is studied via the Euler-Maclaurin expansion. An optimization problem is formulated, and its solution noise schedule is obtained via calculus of variations, shown to follow a tangent law whose coefficient is determined by the eigenvalues of the source covariance matrix. For an alternative scenario, more realistic in practice, where pretrained models have been obtained for some given noise schedules, the KL divergence also provides a measure to compare different time discretization strategies in reverse sampling. Experiments across different datasets and pretrained models demonstrate that the time discretization strategy selected by our approach consistently outperforms baseline and search-based strategies, particularly when the budget on the number of function evaluations is very tight.
@@ I Introduction
Aiming at understanding from first principles, the interplay between noise scheduling and sampling efficiency, in this work, we first utilize the Girsanov theorem and the Cramér-Rao bound to reveal that the KL divergence between the continuous and discretized reverse-time SDEs in the Gaussian setting serves as a strict theoretical lower bound for that of general distributions. This finding validates the Gaussian setting as a necessary proxy for parameter optimization in DMs, providing a strong motivation for leveraging the analytical properties of Gaussians to analyze diffusion dynamics. Consequently, in the remainder of this paper, we adopt an elementary approach focusing on a multivariate Gaussian source with a given covariance matrix. Such a simple model induces a linear form of the optimal posterior estimator, and consequently an explicit closed-form characterization of the evolution trajectory of the distributions across the deterministic reverse sampling steps. Using this closed-form formulation, we can conveniently investigate the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output.
@@
Notation: When there is no ambiguity, we omit the explicit time argument $t$ for time-dependent functions (e.g., writing $f$ for $f(t)$) and denote its discrete-time counterpart $f(t_{j})$ by $f_{t_{j}}$. Derivatives with respect to $t$ are written as $\dot{f}=\mathrm{d}f/\mathrm{d}t$ and $\ddot{f}=\mathrm{d}^{2}f/\mathrm{d}t^{2}$. Throughout this paper, $\log(\cdot)$ represents the natural logarithm.
@@ II Preliminaries
In this section, we first review the mechanism of DMs, encompassing the forward diffusion process with its noise schedule design, and the reverse sampling process employed for generation.
@@
where $t\in[0,1]$ denotes the diffusion time, $\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ represents the source data sample, and $\alpha_{t},\sigma_{t}$ parameterize the noise schedule that directly controls the noise-adding process. Commonly adopted noise schedules in DMs are categorized into two types: variance-preserving (VP) and variance-exploding (VE) [25]. The VP schedule is constrained by $\alpha_{t}^{2}+\sigma_{t}^{2}=1$, with boundary conditions $\lim_{t\to 0}\alpha_{t}=1$ and $\lim_{t\to 1}\alpha_{t}=0$. In contrast, the VE schedule maintains $\alpha_{t}\equiv 1$ while satisfying $\lim_{t\to 0}\sigma_{t}=0$ and $\lim_{t\to 1}\sigma_{t}=\sigma_{\max}\gg 1$. Despite these differences, the forward transition kernel for both schedules unifies to:
$q_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ (2)
which progressively diffuses the source distribution $q_{0}(\mathbf{x}_{0})$ toward a tractable Gaussian prior as $t\to 1$. Moreover, the trajectory of the marginal distributions governed by (2) can be equivalently described by a SDE [25, 11]:
$\mathrm{d}\mathbf{x}_{t}=f(t)\mathbf{x}_{t}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (3)
where $\mathbf{w}_{t}$ denotes the standard Wiener process. The drift coefficient $f(t)$ and diffusion coefficient $g(t)$ are derived as [11]:
$f(t)=\frac{\dot{\alpha}}{\alpha},\quad g(t)=\sqrt{2\sigma\left(\dot{\sigma}-\frac{\dot{\alpha}}{\alpha}\sigma\right)}.$ (4)
According to the classical theory of diffusion processes [1, 25], the corresponding reverse-time SDE flowing from $t=1$ to 0 is given by:
$\mathrm{d}\mathbf{x}_{t}=\left(f(t)\mathbf{x}_{t}-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})\right)\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}_{t},$ (5)
where $\bar{\mathbf{w}}_{t}$ is a standard Wiener process in the reverse time. Theoretically, when initialized at $\mathbf{x}_{1}\sim q_{1}(\mathbf{x}_{1})$, this process reverses the diffusion trajectory and enables sampling from the source distribution $q_{0}(\mathbf{x}_{0})$.
To render the reverse-time SDE tractable, DMs employ a neural network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ to approximate the scaled score function, i.e., $\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$, which is optimized via the following mean squared error (MSE) objective:
@@
$\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}\left[\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}}\bigg|\mathbf{x}_{t}\right]=\frac{\mathbf{x}_{t}-\alpha_{t}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\sigma_{t}}=-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}),$ (7)
@@
where the last equality follows from Tweedie’s formula [4]. Thus, learning to predict the noise $\bm{\epsilon}$ is mathematically equivalent to estimating the posterior mean of the source data $\mathbf{x}_{0}$.
To derive an implementable numerical scheme, we discretize the reverse-time SDE (5). Consider a time discretization sequence $\{t_{i}\}_{i=0}^{N}$ decreasing from $t_{N}=1$ to $t_{0}=0$. For the integration step from $t_{j}$ to $t_{j-1}$ (i.e., the interval $t\in(t_{j-1},t_{j}]$), we approximate the score component by freezing the neural estimator at the starting time $t_{j}$. This yields the following piecewise SDE approximation [2, 15]:
@@
$\mathrm{d}\mathbf{x}_{t}=\left(f(t)\mathbf{x}_{t}-h(t)\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}),$ (8)
where the coefficient $h(t)$ is derived as:
$h(t)=\frac{g^{2}(t)}{-\sigma(t)}=2\left(\frac{\dot{\alpha}}{\alpha}\sigma-\dot{\sigma}\right).$ (9)
In practice, the initialization parameter $\sigma_{N}$ is set by the noise schedule, with $\sigma_{N}=1$ for the VP setting and $\sigma_{N}=\sigma_{\max}$ for the VE setting. For subsequent theoretical analysis, we denote the path measures induced by the exact reverse-time SDE (5) and the approximate reverse-time SDE (8) as $\mathbb{Q}$ and $\mathbb{P}$, respectively.
While the stochastic sampling above recovers the data distribution, deterministic sampling is often preferred in practice for its stability and effectiveness. By converting the SDE into its corresponding probability flow ODE (PF-ODE) [25], we obtain the deterministic reverse sampling process [24, 16]. This process inverts the forward perturbation to generate samples from $q_{0}(\mathbf{x}_{0})$ via the following iterative update rule for $j=N,N-1,\dots,1$:
where the process is initialized with $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})$.
III Gaussian Reference and Lower Bound
This section provides the theoretical motivation for our focus on the Gaussian setting. Specifically, by utilizing stochastic analysis on the reverse-time SDEs, we demonstrate that the Gaussian case serves as a fundamental lower bound for general distributions.
We first introduce a Gaussian reference process. Let $q_{0,G}(\mathbf{x}_{0})$ be a Gaussian distribution constructed to share the identical covariance structure with the underlying data distribution $q_{0}(\mathbf{x})$. Specifically, denoting the data covariance as $\mathrm{Cov}_{q_{0}}[\mathbf{x}_{0}]=\bm{\Sigma}_{\mathbf{x}}$, we consider the reference distribution $q_{0,G}$ with covariance $\bm{\Sigma}_{\mathbf{x}}$. Due to the linearity of the forward diffusion governed by (1), this covariance matching is preserved throughout the forward process, i.e. $\mathrm{Cov}_{q_{t}}[\mathbf{x}_{t}]=\mathrm{Cov}_{q_{t,G}}[\mathbf{x}_{t}]=\bm{\Sigma}_{t}=\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}$. Moreover, the Hessian of the log-density for this Gaussian reference is given directly by:
$\nabla_{\mathbf{x}_{t}}^{2}\log q_{t,G}(\mathbf{x}_{t})=-\bm{\Sigma}_{t}^{-1}.$ (11)
Analogous to the general case, we define $\mathbb{Q}_{G}$ and $\mathbb{P}_{G}$ as the path measures induced by the exact reverse-time SDE and the approximate reverse-time SDE, respectively, operating under the Gaussian source assumption $q_{0,G}$.
Theorem 1.
Under the optimal estimation, the KL divergence of the Gaussian reference process serves as a lower bound for the true process:
$D_{KL}(\mathbb{Q}\|\mathbb{P})\geq D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})\geq D_{KL}(q_{0,G}\|p_{0,G}).$ (12)
Proof.
We first aim to prove the inequality $D_{KL}(\mathbb{Q}\|\mathbb{P})\geq D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})$. Let $\Delta$ denote the gap between the KL divergence of the general process and that of the Gaussian reference process.
Under the assumption that the neural network approximates the optimal score, i.e., $\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ given in (7), its Jacobian is given by $\nabla_{\mathbf{x}}\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}_{t}}^{2}\log q_{t}(\mathbf{x}_{t})$. Specifically for the Gaussian reference process, this simplifies to a constant Jacobian $\nabla_{\mathbf{x}}\bm{\epsilon}^{*}_{\theta,G}(\mathbf{x}_{t},t)=\sigma_{t}\bm{\Sigma}_{t}^{-1}$ due to (11). Consequently, using Lemma 3 in Appendix E, the difference decomposes as:
$\Delta$ $\triangleq D_{KL}(\mathbb{Q}\|\mathbb{P})-D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})$
$=\Delta_{init}+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\left(\mathbb{E}_{q_{t_{j}}}\left[\|-\sigma_{t_{j}}\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\sigma_{t_{j}}\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)+\mathcal{R}(\delta),$ (13)
where the initial term $\Delta_{init}$ is defined as:
$\Delta_{init}=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))-D_{KL}(q_{1,G}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})).$ (14)
For the initial term $\Delta_{\text{init}}$, we utilize the decomposition of KL divergence: $D_{KL}(q\|p)=-H(q)+H(q,p)$, where $H(q,p)=-\mathbb{E}_{q}[\log p]$ is the cross-entropy. Since the prior $p_{\text{prior}}=\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})$ is Gaussian, the cross-entropy term depends solely on the first two moments of $q$. By construction, the Gaussian reference $q_{1,G}$ shares the identical mean and covariance with the true distribution $q_{1}$, we have
$\mathbb{E}_{q_{1}}[\log p_{\text{prior}}]=\mathbb{E}_{q_{1,G}}[\log p_{\text{prior}}].$ (15)
The difference $\Delta_{\text{init}}$ thus simplifies to the difference in differential entropies
$\Delta_{\text{init}}=\left(-H(q_{1})\right)-\left(-H(q_{1,G})\right)=H(q_{1,G})-H(q_{1}).$ (16)
According to the Maximum Entropy Theorem [3], among all distributions with a fixed covariance matrix, the Gaussian distribution maximizes the differential entropy. Therefore, $H(q_{1,G})\geq H(q_{1})$, implying $\Delta_{\text{init}}\geq 0$.
To determine the sign of $\Delta$, we analyze the Jacobian terms. The difference can be expanded as:
$\mathbb{E}_{q_{t_{j}}}\left[\|-\sigma_{t_{j}}\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\sigma_{t_{j}}\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}$
$=\sigma_{t_{j}}^{2}\left(\mathbb{E}_{q_{t_{j}}}\left[\|\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (17)
$\geq\sigma_{t_{j}}^{2}\left(\left\|\mathbb{E}_{q_{t_{j}}}[\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})]\right\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (18)
$=\sigma_{t_{j}}^{2}\left(\|-\mathcal{I}(q_{t_{j}})\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (19)
$=\sigma_{t_{j}}^{2}\left(\|\mathcal{I}(q_{t_{j}})\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$
$\geq 0,$ (20)
where $\mathcal{I}(q_{t})\triangleq\mathbb{E}_{q_{t}}[-\nabla_{\mathbf{x}}^{2}\log q_{t}(\mathbf{x})]$ denotes the Fisher Information Matrix. The inequality in (18) follows from Jensen’s inequality applied to the convex function $\|\cdot\|_{F}^{2}$. Equality (19) utilizes the identity that the expected Hessian of the log-likelihood is the negative of the Fisher Information Matrix. Finally, the inequality in (20) is guaranteed by the multivariate Cramér-Rao Bound, which states that $\mathcal{I}(q_{t})\succeq\bm{\Sigma}_{t}^{-1}$ [3], and all the equalities hold if and only if $q_{t}(\mathbf{x})$ is also a Gaussian distribution.
Substituting this result and $\Delta_{init}\geq 0$ into (III), we obtain $\Delta\geq 0$ for a sufficiently small step size. Furthermore, the second inequality $D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})\geq D_{KL}(q_{0,G}\|p_{0,G})$ follows directly from the Data Processing Inequality (DPI), as the marginal distribution at $t=0$ is a projection of the path measure. ∎
Theorem 1 reveals that the Gaussian setting characterizes a theoretical lower bound of the KL divergence for general distributions. This motivates our subsequent analysis of the Gaussian source, serving as a tractable proxy for general distributions, to derive optimal parameter designs for DMs.
IV KL-Based Analysis on Gaussian Setting and Applications
In this section, we present the results derived from our theoretical analysis under the Gaussian setting and the widely used deterministic reverse sampling process introduced in (10), and demonstrate their applications. For clarity of exposition, we focus on the VP setting; analogous results for the VE setting are briefly described in discussions and remarks.
As the setup given in Section III, we suppose $\mathbf{x}_{0}\sim q_{0,G}(\mathbf{x}_{0})=\mathcal{N}(\mathbf{0},\bm{\Sigma}_{\mathbf{x}})$, so the marginal distribution $q_{t}(\mathbf{x}_{t})$ remains Gaussian:
$q_{t,G}(\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{0},\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right).$ (21)
@@ IV KL-Based Analysis on Gaussian Setting and Applications
Moreover, in the Gaussian setting, the optimal posterior estimator admits a closed-form solution [20]:
@@
Without loss of generality, we assume that the covariance matrix $\bm{\Sigma}_{\mathbf{x}}$ is positive definite. Let the eigendecomposition of the covariance matrix be $\bm{\Sigma}_{\mathbf{x}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{-1}$, where $\mathbf{U}$ is an orthogonal matrix and $\mathbf{\Lambda}=\text{diag}(\mu_{1},\ldots,\mu_{k})$ denotes the diagonal matrix of eigenvalues.
IV-A KL Divergence Analysis